# RESSA: Repair Sparse Vision-Language Models via Sparse Cross-Modality Adaptation

### v1

@@ 7 Conclusion
In this paper, motivated by the challenges associated with deploying VLMs in real-world applications, we investigate the potential of pruning VLMs. Specifically, recognizing that VLMs encompass models from different modalities, we conduct empirical studies to explore the distribution of sparsity ratios across these models and how sparsity impacts performance, thereby highlighting the necessity of repairing pruned VLMs. Subsequently, we introduce RESSA, which addresses this challenge by repairing pruned VLMs through cross-modality adaptation and SparseLoRA finetuning. Extensive experiments validate the effectiveness of RESSA, providing valuable insights for future research on VLM sparsity.

### v2

@@ 7 Conclusion
In this paper, motivated by the challenges associated with deploying VLMs in real-world applications, we investigate the potential of pruning VLMs. Specifically, recognizing that VLMs encompass models from different modalities, we conduct empirical studies to explore the distribution of sparsity ratios across these models and how sparsity impacts performance, thereby highlighting the necessity of restoring pruned VLMs. Subsequently, we introduce MAF, which addresses this challenge by restoring pruned VLMs through cross-modality adaptation and SparseLoRA finetuning. Extensive experiments validate the effectiveness of MAF, providing valuable insights for future research on VLM sparsity.
8 Limitations
Despite our progress, limitations remain in our work. Although our proposed methods are universal for all VLM models, we have primarily focused on BLIP family models and selected tasks . We believe our methods can be easily extended to a broader range of models and tasks. On the other hand, given there may be potentially high-quality dataset for restoring pruned models, we believe the incorporation of such datasets would further promotes our proposed methods